OpenStudy (anonymous):

Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this.

OpenStudy (anonymous):

What's the question?

OpenStudy (anonymous):

The function(s)

OpenStudy (anonymous):

x + 8y = 14 -3x + 4y = 0 Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this.